Show that if a and b are :
(a). in the same direction then |a + b| = |a| +|b|.
(b). in the opposite direction then |a − b| = |a| + |b|.
1
Expert's answer
2019-10-04T10:11:39-0400

Align both vectors along the X-axis of cartesian coordinates. Let the vector "a" be directed in a positive direction. Then we can write:

"a=(a_x,0,0,\dots,0)\\"

"b=(b_x,0,0,\dots,0)\\"


If vectors are in the opposite direction, then "|a| = a_x>0, |b| = -b_x >0.\\"

"|a-b| = |(a_x-b_x,0,0,\dots,0)|=\sqrt{(a_x-b_x)^2}=\\"

"a_x-b_x=|a|+|b|.\\"


If vectors are in the same direction, then "|a| = a_x>0, |b| = b_x >0.\\"

"|a+b| = |(a_x+b_x,0,0,\dots,0)|=\sqrt{(a_x+b_x)^2}=\\"

"a_x+b_x=|a|+|b|.\\"